OpenStudy (anonymous):

(6x+1)^2 +8(6x+1) + 16=0

OpenStudy (anonymous):

let x be 6x+1 x^2+8x+16=0 (x+4)(x+4)=0 sub back in (6x+1+4)=0 6x=-5 x=-5/6